Unlined channel to be excavated in stiff clay to convey 9 m3/s. The slope is 0.0028. Design the dimensions.
Answer #1

2- 0 62Now Area b= δ9129-3.606 = 3.566m 0.6657m n y= 3.566, b=0 9129-3606 x 3566b= 0.9129-3.606 x 0.6657 b= 6.SI2ym Rahel tu uidAng channe y 6.5% 67 m